OpenStudy (anonymous):

find the vertex the line of symmetry the maximum or minimum value of the quadratic function f(x)=3x^2-18+13

OpenStudy (anonymous):

vertex at \[\frac{b}{2a}=-\frac{-18}{2\times 3}=3\] for the first coordinate. find the second by replacing x by 3 to see what you get

OpenStudy (anonymous):

i am assuming you meant \[f(x)=3x^2-18x+13\]

OpenStudy (anonymous):

you should get \[f(3)=-14\] so vertex is (3,-14) and that is a minimum since parabola opens up (leading coefficient is positive)